# Distributed Differentially Private Data Analytics via Secure Sketching

### v1

@@ Appendix C Utility Proofs
Denote by $\mathbf{G}_{i}$ the $i$th row of $\mathbf{G}$. We start by observing that $\mathbf{G}_{i}^{T}\mathbf{x}$ is Gaussian distributed with mean 00 and variance $\sigma^{2}\cdot\|\mathbf{x}\|^{2}$. Thus, we focus our attention on controlling $\frac{\|\mathbf{G}\mathbf{x}\|^{2}}{\sigma^{2}\cdot\|\mathbf{x}\|^{2}}$, that is, we assume that $\mathbf{x}$ is a unit vector and that the entries of $\mathbf{G}$ are standard normal Gaussian random variables. The overall claim then follows by rescaling. Consider an $\varepsilon$-net $N_{\varepsilon}$ of $V$, that is, for every $\mathbf{x}\in V$ with unit norm there exists a vector $\mathbf{x}^{\prime}\in N_{\gamma}$ with $\|\mathbf{x}-\mathbf{x}^{\prime}\|\leq\varepsilon$. Such nets exist with $|N_{\gamma}|\leq\exp(\eta\cdot k\log 1/\gamma)$ for some absolute constant $\eta$, see [56]. Suppose $\gamma=1/4$. Using concentration bounds for sums of Gaussians (see for example Lemma 1 of [48]), we then have for any $\|x^{\prime}\|$ with probability $1-\beta$
@@
We now extend this argument to all vectors, using an argument from [3] (Lemma 4 of that reference). Let $\mathbf{U}$ be an orthogonal basis of $V$. Then our goal is to control $\|\mathbf{G}\mathbf{x}\|^{2}=\mathbf{x}^{T}\mathbf{U}^{T}\mathbf{G}^{T}\mathbf{G}\mathbf{U}x$. Define the matrix $\mathbf{B}:=\mathbf{U}^{T}\mathbf{G}^{T}\mathbf{G}\mathbf{U}-n\cdot\mathbf{I}$, where $\mathbf{I}$ is the identity matrix. Note that $\mathbb{E}[\mathbf{U}^{T}\mathbf{G}^{T}\mathbf{G}\mathbf{U}]=n\cdot\mathbf{I}$ and that $\|\mathbf{B}\mathbf{x}\|$ is the deviation of $\|\mathbf{G}\mathbf{x}\|$ around its expectation. Let $\|B\|_{op}:=\sup_{\mathbf{x}\in V}\|Bx\|$.
@@
Proof of Lemma 4.1.
We first argue that $\mathbf{G}$ is Gaussian distributed. Each $G_{i}$ has independent Gaussian entries and multiplying a Gaussian with a random Rademacher does not change the distribution. Therefore, the entries of $S_{i}G_{i}$ are likewise Gaussian distributed, with mean 00 and variance at most $(1+\gamma)\frac{n}{m}\sigma^{2}\leq 2\frac{n}{m}\sigma^{2}$ due to Lemma B.1. Concluding, the variance of the entries of $\mathbf{G}=\sum_{i\in[s]}S_{i}G_{i}$ is therefore at most $2s\frac{n}{m}\sigma^{2}$. Applying Lemma C.2, we then have with probability $1-\beta$
$\sup_{\mathbf{x}\in V}\|\mathbf{G}\mathbf{x}\|_{2}^{2}$ $\leq 2\frac{n}{m}\sigma^{2}\cdot\|\mathbf{x}\|^{2}$
$+\eta\cdot 2s\frac{n}{m}\sigma^{2}\cdot\|\mathbf{x}\|^{2}\cdot(\sqrt{(k+\log 1/\beta)\cdot m}+(k+\log 1/\beta)).$
@@
Proof of Theorem 4.4.
@@
Let $\varepsilon$, $\delta$, $R_{\sigma}$ and $T_{\mathbf{S}}$ be given as in the theorem. We consider the output $A_{reg}(T_{\mathbf{S}}(R_{\sigma}(\mathbf{a}_{1}),\dots,R_{\sigma}(\mathbf{a}_{n})))$.
@@
If $n\leq 8m\ln(dm/\delta)+t^{\prime}$, then the $R_{\sigma}$ outputs $0^{d}$, then the optimal $\mathbf{x}^{\prime}=0^{d}$, leading to error $\lVert\mathbf{A}\mathbf{x}^{\prime}-\mathbf{b}\rVert_{2}^{2}+\lambda\lVert\mathbf{x}^{\prime}\rVert_{2}^{2}=\lVert\mathbf{b}\rVert^{2}\leq\eta^{\prime}n\leq\eta^{\prime}m\ln(dm/\delta)+t^{\prime}$.
@@
Otherwise, $\mathcal{M}(\mathbf{A})=T_{\mathbf{S}}(R_{\sigma}(\mathbf{a}_{1}),\dots,R_{\sigma}(\mathbf{a}_{n}))=\frac{1}{\sqrt{s}}\sum_{i\in[s]}\mathbf{S}_{i}(\mathbf{A}+\mathbf{G}_{i})$ where $\mathbf{a}_{i}$ denotes row $i$ of $\mathbf{A}\in\mathbb{R}^{n\times d}$ and $\mathbf{G}\leftarrow\mathcal{N}(0,\sigma^{2})^{n\times d}$, is $(\varepsilon,\delta)$ differentially private.
@@
for either $\mathbf{x}=\mathbf{x}_{\text{OPT}}$ or $\mathbf{x}=\mathbf{x}^{\prime}$ and for a sufficiently large constant $\eta.$ Then we have
$\|\mathbf{A}\mathbf{x}^{\prime}-\mathbf{b}\|^{2}+\lambda\|x^{\prime}\|^{2}\leq(1+\alpha_{S})\|\mathbf{S}(\mathbf{A}\mathbf{x}^{\prime}-\mathbf{b})\|^{2}+\lambda\|x^{\prime}\|^{2}$
@@
$=$ $(1+\alpha_{S})\frac{1}{s}\|\sum_{i\in[s]}\mathbf{S}_{i}((\mathbf{A}+\mathbf{G}_{i}-\mathbf{G}_{i})\mathbf{x}^{\prime}-\mathbf{b}+\mathbf{g}_{i}-\mathbf{g}_{i})\|^{2}+\lambda\|x^{\prime}\|^{2}$
@@
$\leq$ $(1+\alpha_{S})^{2}\cdot\left(\frac{1}{s}\|\sum_{i\in[s]}S_{i}((\mathbf{A}+\mathbf{G}_{i})\mathbf{x}^{\prime}-(\mathbf{b}+\mathbf{g}_{i})\|^{2}+\lambda\cdot\|\mathbf{x}^{\prime}\|^{2}\right)$
(2) $+2\cdot\left(1+\frac{1}{\alpha_{S}}\right)\eta\sigma^{2}nd\log\frac{1}{\beta}(\|\mathbf{x}^{\prime}\|^{2}+1)$
@@
$\|\mathbf{x}^{\prime}\|\leq\frac{1}{\lambda}\cdot\left(\frac{1}{s}\|\sum_{i\in[s]}S_{i}((\mathbf{A}+\mathbf{G}_{i})\mathbf{x}_{\text{OPT}}-(\mathbf{b}+\mathbf{g}_{i}))\|^{2}+\lambda\cdot\|\mathbf{x}_{\text{OPT}}\|^{2}\right).$

### v2

@@ 5. Numerical Linear Algebra in LTM
Denote by $\mathbf{G}_{i}$ the $i$th row of $\mathbf{G}$. We start by observing that $\mathbf{G}_{i}^{T}\mathbf{x}$ is Gaussian distributed with mean 00 and variance $\sigma^{2}\cdot\|\mathbf{x}\|^{2}$. Thus, we focus our attention on controlling $\frac{\|\mathbf{G}\mathbf{x}\|^{2}}{\sigma^{2}\cdot\|\mathbf{x}\|^{2}}$, that is, we assume that $\mathbf{x}$ is a unit vector and that the entries of $\mathbf{G}$ are standard normal Gaussian random variables. The overall claim then follows by rescaling. Consider an $\varepsilon$-net $N_{\varepsilon}$ of $V$, that is, for every $\mathbf{x}\in V$ with unit norm there exists a vector $\mathbf{x}^{\prime}\in N_{\gamma}$ with $\|\mathbf{x}-\mathbf{x}^{\prime}\|\leq\varepsilon$. Such nets exist with $|N_{\gamma}|\leq\exp(\eta\cdot k\log 1/\gamma)$ for some absolute constant $\eta$, see (Pisier, 1999). Suppose $\gamma=1/4$. Using concentration bounds for sums of Gaussians (see for example Lemma 1 of (Laurent and Massart, 2000)), we then have for any $\|x^{\prime}\|$ with probability $1-\beta$
@@
We now extend this argument to all vectors, using an argument from (Arora et al., 2006) (Lemma 4 of that reference). Let $\mathbf{U}$ be an orthogonal basis of $V$. Then our goal is to control $\|\mathbf{G}\mathbf{x}\|^{2}=\mathbf{x}^{T}\mathbf{U}^{T}\mathbf{G}^{T}\mathbf{G}\mathbf{U}x$. Define the matrix $\mathbf{B}:=\mathbf{U}^{T}\mathbf{G}^{T}\mathbf{G}\mathbf{U}-n\cdot\mathbf{I}$, where $\mathbf{I}$ is the identity matrix. Note that $\mathbb{E}[\mathbf{U}^{T}\mathbf{G}^{T}\mathbf{G}\mathbf{U}]=n\cdot\mathbf{I}$ and that $\|\mathbf{B}\mathbf{x}\|$ is the deviation of $\|\mathbf{G}\mathbf{x}\|$ around its expectation. Let $\|B\|_{op}:=\sup_{\mathbf{x}\in V}\|Bx\|$.
@@
We will use the following spectral bounds for both linear regression and low rank approximation. We believe that there may be further applications and that the bounds themselves are therefore of independent interest.
Lemma 5.2.
Let $\mathbf{S}\in\frac{1}{\sqrt{s}}\cdot\{-1,0,1\}^{m\times n}$ be an $(\alpha,\beta,m,s)$ OSNAP with $\mathbf{S}=\frac{1}{\sqrt{s}}\sum_{i\in[s]}\mathbf{S}_{i}$, where $\mathbf{S}_{i}\in\{-1,0,1\}^{n\times m}$ has exactly one non-zero entry per row. Let $\mathbf{G}=\sum_{i\in[s]}\mathbf{S}_{i}\mathbf{G}_{i}$ where every matrix $\mathbf{G}_{i}\in\mathbb{R}^{n\times d}$ has independent Gaussian entries $\mathcal{N}(0,\sigma^{2})$. Further, let $V$ be a set of $d$-dimensional vectors lying in a $k$-dimensional subspace. Then with probability at least $1-\beta$ for some absolute constant $\eta$
$\sup_{\mathbf{x}\in V}\frac{1}{s}\|\mathbf{G}\mathbf{x}\|_{2}^{2}\leq\eta\cdot 2\frac{n}{m}\sigma^{2}\cdot\|\mathbf{x}\|^{2}\cdot(\sqrt{(k+\log 1/\beta)\cdot m}$
$\hskip 113.81102pt+(k+\log 1/\beta)).$
and
Proof.
We first argue that $\mathbf{G}$ is Gaussian distributed. Each $G_{i}$ has independent Gaussian entries and multiplying a Gaussian with a random Rademacher does not change the distribution. Therefore, the entries of $S_{i}G_{i}$ are likewise Gaussian distributed, with mean 00 and variance at most $(1+\gamma)\frac{n}{m}\sigma^{2}\leq 2\frac{n}{m}\sigma^{2}$ due to Lemma 4.5. Concluding, the variance of the entries of $\mathbf{G}=\sum_{i\in[s]}S_{i}G_{i}$ is therefore at most $2s\frac{n}{m}\sigma^{2}$. Applying Lemma 5.1, we then have with probability $1-\beta$
$\sup_{\mathbf{x}\in V}\|\mathbf{G}\mathbf{x}\|_{2}^{2}\leq 2\frac{n}{m}\sigma^{2}\cdot\|\mathbf{x}\|^{2}+\eta\cdot 2s\frac{n}{m}\sigma^{2}\cdot\|\mathbf{x}\|^{2}$
$\cdot(\sqrt{(k+\log 1/\beta)\cdot m}+(k+\log 1/\beta)).$
@@
Related, though slightly weaker bounds can also be derived for the Gamma/Laplace mechanism. While it is possible to derive bounds for sparse Rademacher sketches, we focus on the dense constructions as these yield $\varepsilon$ differentially private mechanisms (as opposed to $(\varepsilon,\delta)$ differentially private mechanisms).
Lemma 5.3.
Let $\mathbf{S}\in\frac{1}{\sqrt{s}}\cdot\{-1,0,1\}^{m\times n}$ be an $(\alpha,\beta,m,m)$ OSNAP with $\mathbf{S}=\frac{1}{\sqrt{m}}\sum_{i\in[s]}\mathbf{S}_{i}$, where $\mathbf{S}_{i}\in\{-1,0,1\}^{m\times n}$ has exactly one non-zero entry per row and $n/m$ non-zero entries per column. Let $\mathbf{L}=\frac{1}{\sqrt{m}}\sum_{i\in[m]}\mathbf{S}_{i}\mathbf{\Gamma}_{i}$ where every matrix $\mathbf{\Gamma}_{i}\in\mathbb{R}^{n\times d}$ has independent entries $X_{i}-Y_{i}$, where $X_{i},Y_{i}\sim\Gamma(m/n,b)$. Then with probability at least $1-\beta$ for some absolute constant $\eta$ and
$\|\mathbf{L}\|_{F}^{2}\leq\eta dm^{2}b^{2}\log^{2}\frac{m}{\beta}$
Proof.
We have $\|\mathbf{L}\|_{F}\leq\sqrt{m}\cdot\|\max_{i\in[m]}\mathbf{S}_{i}\mathbf{\Gamma}_{i}\|_{F}$, thus it is sufficient to give a high probability bound on $\|\mathbf{S}_{i}\mathbf{\Gamma}_{i}\|_{F}$.
The entries of $\mathbf{S}_{i}\mathbf{\Gamma}_{i}$ are Laplace distributed entries with mean 00 and scale parameter $b$ by infinite divisibility of a Laplace random variable into $\Gamma$ distributed random variables. Thus the expected value of $\|\mathbf{S}_{i}\mathbf{\Gamma}_{i}\|_{F}^{2}$ is $md\cdot 2b^{2}$.
To prove high concentration, let us consider the moments of the Laplace distribution. For a Laplace random variable $L$ with mean 00 and scale $b$, we have $\mathbb{E}[L^{2\ell}]=(2\ell)!b^{2\ell}$. Then due to Hoelder’s inequality
$\mathbb{E}\left[\|\mathbf{S}_{i}\mathbf{\Gamma}_{i}\|_{F}^{2\ell}\right]=$ $\mathbb{E}\left[\left(\sum_{j=1}^{md}L_{j}^{2}\right)^{\ell}\right]$
$\leq$ $\mathbb{E}\left[\left(\sum_{j=1}^{md}L_{i}^{2\ell}\right)\right]\cdot\left(\sum_{j=1}^{md}1^{\frac{\ell}{\ell-1}}\right)^{\ell-1}$
$=$ $(2\ell)!b^{2\ell}\cdot(md)^{\ell}$
Applying Markov’s inequality, we then have for any $\gamma>1$
$\mathbb{P}\left[\|\mathbf{S}_{i}\mathbf{\Gamma}_{i}\|_{F}^{2}>\gamma\cdot md\cdot 2b^{2}\right]$
$=$ $\mathbb{P}\left[\|\mathbf{S}_{i}\mathbf{\Gamma}_{i}\|_{F}^{2\ell}>\left(\gamma\cdot md\cdot 2b^{2}\right)^{\ell}\right]$
$\leq$ $\frac{(2\ell)!b^{2\ell}(md)^{\ell}}{\left(\gamma\cdot md\cdot 2b^{2}\right)^{\ell}}=\frac{(2\ell)!}{(2\gamma)^{\ell}}$
$\leq$ $\frac{(2\ell)^{2\ell}}{(2\gamma e^{2})^{\ell}}\sqrt{6\pi\ell}$
For $\gamma>4\ell^{2}$ and $\ell=\log\frac{m}{\beta}$, this term is at most $\beta/m$. We can therefore conclude that with probability $1-\beta$, $\max_{i\in[m]}\mathbf{S}_{i}\mathbf{\Gamma}_{i}\leq 4\log\frac{m}{\beta}$ and thus $\|\mathbf{L}\|_{F}^{2}\leq\eta m^{2}db^{2}\log^{2}\frac{m}{\beta}$, for some absolute constant $\eta$. ∎
We first begin with our utility results for low rank approximation. Using the Gaussian mechanism, we obtain the following guarantee.
Theorem 5.4.
Let $\varepsilon\geq 0$, $\delta\in(0,1)$, $t^{\prime}<n/2$. Let $\mathbf{S}\sim\mathcal{D}_{m,n}^{\mathtt{sketch}}$. Define $A_{k}$ as an algorithm that computes $\text{argmin}_{\mathbf{X}\text{ rank }k,\in\mathbb{R}^{n\times d}}\lVert\mathbf{A}-\mathbf{A}\mathbf{X}\mathbf{X}^{T}\rVert_{F}^{2}$. An instantiation of the trusted computation model $(R_{D},T_{\mathbf{S}},A_{k})$ specified in Equations 2 and 3, using noise distribution $D\sim\mathcal{N}(0,\sigma^{2})$ according to Corollary 4.7, computes an orthogonal projection $\mathbf{X}^{\prime}$ that with probability $1-\beta$ achieves a multiplicative error $(1+O(\alpha_{\mathbf{S}}))$ and additive error $O\left(\frac{kd^{3/2}}{\alpha_{\mathbf{S}}^{3}}\log^{3/2}(1/\beta)\varepsilon^{-1}\log^{1/2}(1/\delta)\right)$ for the low rank approximation problem.
Proof.
Lemma 5.5 (Generalized Triangle Inequality (Becchetti et al., 2019)).
For any two real numbers $a,b$ and any $\alpha>0$
Finally, we require the following consequence of subspace embedding properties.
Lemma 5.6 (Paraphrased from Definition 1, Lemma 11, and Theorem 12 (Cohen et al., 2015)).
$\mathbf{S}$ be drawn from an OSNAP distribution. Then for any matrix $A\in\mathbb{R}^{n\times d}$ and any rank $k$ orthogonal matrix $\mathbf{X}\in\mathbb{R}^{d\times k}$, we have
$\|\mathbf{S}\mathbf{A}-\mathbf{S}\mathbf{A}\mathbf{X}\mathbf{X}^{T}\|_{F}=(1\pm\alpha)\cdot\|\mathbf{A}-\mathbf{A}\mathbf{X}\mathbf{X}^{T}\|_{F}.$
Without noise, the low rank approximation outputs $\mathbf{X}_{\text{OPT}}=\text{argmin}_{\mathbf{X}\text{ rank }k,\in\mathbb{R}^{n\times d}}\lVert\mathbf{A}-\mathbf{A}\mathbf{X}\mathbf{X}^{T}\rVert_{F}^{2}$, and after sketching with Gaussian noise of variance $\sigma^{2}$ outputs $\mathbf{X}^{\prime}=\text{argmin}_{\mathbf{X}\text{ rank }k,\in\mathbb{R}^{n\times d}}\lVert(\mathbf{A}+\mathbf{G})-(\mathbf{A}+\mathbf{G})\mathbf{X}\mathbf{X}^{T}\rVert_{F}^{2}$. Let $\mathbf{X}^{\prime}$ be the matrix returned by the mechanism. We have
$\lVert$ $\mathbf{A}-\mathbf{A}\mathbf{X}^{\prime}\mathbf{X}^{\prime T}\rVert_{F}$
$\leq(1+\alpha_{\mathbf{S}})\sqrt{\frac{1}{s}}\left\lVert\sum_{i\in[s]}\mathbf{S}_{i}\mathbf{A}-\mathbf{S}_{i}\mathbf{A}\mathbf{X}^{\prime}\mathbf{X}^{\prime T}\right\rVert_{F}$
$\leq(1+\alpha_{\mathbf{S}})\sqrt{\frac{1}{s}}$
$\left\lVert\sum_{i\in[s]}\mathbf{S}_{i}(\mathbf{A}+\mathbf{G}_{i})-\mathbf{S}_{i}(\mathbf{A}+\mathbf{G}_{i})\mathbf{X}^{\prime}\mathbf{X}^{\prime T}\right\rVert_{F}$
$+(1+\alpha_{\mathbf{S}})\sqrt{\frac{1}{s}}\left\lVert\sum_{i\in[s]}\mathbf{S}_{i}\mathbf{G}_{i}-\mathbf{S}_{i}\mathbf{G}_{i}\mathbf{X}^{\prime}\mathbf{X}^{\prime T}\right\rVert_{F}$
where the first inequality follows from the subspace embedding property. By optimality of $\mathbf{X}^{\prime}$ for the low rank approximation problem on $\sum_{i\in[s]}\mathbf{S}_{i}(\mathbf{A}+\mathbf{G}_{i})$, we then have
$\lVert$ $\mathbf{A}-\mathbf{A}\mathbf{X}^{\prime}\mathbf{X}^{\prime T}\rVert_{F}$
$\leq(1+\alpha_{\mathbf{S}})\sqrt{\frac{1}{s}}$
$~{}~{}\cdot\left\lVert\sum_{i\in[s]}\mathbf{S}_{i}(\mathbf{A}+\mathbf{G}_{i})-\mathbf{S}_{i}(\mathbf{A}+\mathbf{G}_{i})\mathbf{X}_{\text{OPT}}\mathbf{X}_{\text{OPT}}^{T}\right\rVert_{F}$
$~{}~{}+(1+\alpha_{\mathbf{S}})\sqrt{\frac{1}{s}}\left\lVert\sum_{i\in[s]}\mathbf{S}_{i}\mathbf{G}_{i}-\mathbf{S}_{i}\mathbf{G}_{i}\mathbf{X}^{\prime}\mathbf{X}^{\prime T}\right\rVert_{F}$
$\leq(1+\alpha_{\mathbf{S}})^{2}\left\lVert\mathbf{A}-\mathbf{A}\mathbf{X}_{\text{OPT}}\mathbf{X}_{\text{OPT}}^{T}\right\rVert_{F}$
$~{}~{}~{}~{}~{}~{}~{}+(1+\alpha_{\mathbf{S}})\sqrt{\frac{1}{s}}\left\lVert\sum_{i\in[s]}\mathbf{S}_{i}\mathbf{G}_{i}-\mathbf{S}_{i}\mathbf{G}_{i}\mathbf{X}_{\text{OPT}}\mathbf{X}_{\text{OPT}}^{T}\right\rVert_{F}$
$~{}~{}+(1+\alpha_{\mathbf{S}})\sqrt{\frac{1}{s}}\left\lVert\sum_{i\in[s]}\mathbf{S}_{i}\mathbf{G}_{i}-\mathbf{S}_{i}\mathbf{G}_{i}\mathbf{X}^{\prime}\mathbf{X}^{\prime T}\right\rVert_{F}$
where the inequality follows from Lemma 5.6. Now notice that $\mathbb{I}-\mathbf{X}_{\text{OPT}}\mathbf{X}_{\text{OPT}}^{T}$ and $\mathbb{I}-\mathbf{X}^{\prime}\mathbf{X}^{\prime T}$ are orthogonal projections, multiplying by which cannot increase the Frobenius norm of a matrix. Therefore, for $\mathbf{X}=\mathbf{X}^{\prime}$ or $\mathbf{X}=\mathbf{X}_{\text{OPT}}$:
$\sqrt{\frac{1}{s}}\left\lVert\sum_{i\in[s]}\mathbf{S}_{i}\mathbf{G}_{i}-\mathbf{S}_{i}\mathbf{G}_{i}\mathbf{X}\mathbf{X}^{T}\right\rVert_{F}$
$=\sqrt{\frac{1}{s}}\left\lVert\sum_{i\in[s]}\mathbf{S}_{i}\mathbf{G}_{i}(\mathbb{I}-\mathbf{X}\mathbf{X}^{T})\right\rVert_{F}\leq\sqrt{\frac{1}{s}}\left\lVert\sum_{i\in[s]}\mathbf{S}_{i}\mathbf{G}_{i}\right\rVert_{F}$
Then:
$\lVert\mathbf{A}-\mathbf{A}\mathbf{X}^{\prime}\mathbf{X}^{\prime T}\rVert_{F}$
$\leq(1+\alpha_{\mathbf{S}})^{2}\lVert\mathbf{A}-\mathbf{A}\mathbf{X}_{\text{OPT}}\mathbf{X}_{\text{OPT}}^{T}\rVert_{F}$
(4) $+(1+\alpha_{\mathbf{S}})2\frac{1}{s}\left\lVert\sum_{i\in[s]}\mathbf{S}_{i}\mathbf{G}_{i}\right\rVert_{F}$
Using Lemma 5.2, we have $\frac{1}{s}\left\lVert\sum_{i\in[s]}\mathbf{S}_{i}\mathbf{G}_{i}\right\rVert_{F}^{2}\leq 2n\cdot\sigma^{2}\cdot d\log(1/\beta)$. To simplify the error, observe that for the sketching matrix by (Cohen, 2016) (see Table 2), we can set $s=\frac{\log k}{\alpha_{\mathbf{S}}}$ and $m=\frac{k\log k}{\alpha_{\mathbf{S}}^{2}}\log(1/\beta)$. Using the bound on the variance from Theorem 4.7 and plugging this in above, we can simplify the additive and multiplicative errors.
$\leq(1+\alpha_{\mathbf{S}})^{2}\lVert\mathbf{A}-\mathbf{A}\mathbf{X}_{\text{OPT}}\mathbf{X}_{\text{OPT}}^{T}\rVert_{F}^{2}$
$\phantom{abddddd}+\eta(1+\alpha_{\mathbf{S}})\sqrt{\sigma^{2}\cdot n\cdot d\cdot\log\frac{1}{\tilde{\beta}}}$
$\leq(1+3\alpha_{\mathbf{S}})\lVert\mathbf{A}-\mathbf{A}\mathbf{X}_{\text{OPT}}\mathbf{X}_{\text{OPT}}^{T}\rVert_{F}^{2}$
$\phantom{abddddd}+\tilde{O}\left(\frac{kd^{3/2}}{\alpha_{\mathbf{S}}^{3}}\log^{3/2}(1/\beta)\varepsilon^{-1}\log^{1/2}(1/\delta)\right)$
The stated bound follows by rescaling $\alpha_{\mathbf{S}}$ by a factor of $3$. ∎
Using dense Rademacher sketches and the Gamma/Laplace mechanism, we instead obtain the following guarantee.
Theorem 5.7.
Let $\varepsilon\geq 0$, $\delta\in(0,1)$, $t^{\prime}<n/2$. Let $\mathbf{S}\sim\mathcal{D}_{m,n}^{\mathtt{sketch}}$. Define $A_{k}$ as an algorithm that computes $\text{argmin}_{\mathbf{X}\text{ rank }k,\in\mathbb{R}^{n\times d}}\lVert\mathbf{A}-\mathbf{A}\mathbf{X}\mathbf{X}^{T}\rVert_{F}^{2}$. An instantiation of the trusted computation model
$(R_{D}^{\text{dense}},T_{\mathbf{S}}^{\text{dense}},A_{k})$ specified in Equations 1 and 3, using noise distribution $D$ as the difference between two independent $\Gamma$ random variables parameterized by Corollary 4.10, computes an orthogonal projection $\mathbf{X}^{\prime}$ that with probability $1-\beta$ with multiplicative error $(1+O(\alpha_{\mathbf{S}}))$ and additive error $O(\frac{k^{3}d^{3/2}}{\alpha_{\mathbf{S}}^{6}}\varepsilon^{-1}\log\frac{m}{\beta})$.
Proof.
The proof follows the same line of reasoning as Theorem 5.4. The main difference is that upon reaching Equation 4, we must bound $\frac{1}{m}\left\lVert\sum_{i\in[m]}\mathbf{S}_{i}\mathbf{G}_{i}\right\rVert_{F}^{2}$ via Lemma 5.3 rather than using Lemma 5.2. Assuming the semi-trusted corruption model with at most $t$ corruptions, each entry of $\mathbf{G}_{i}$ consists of $X_{j}-Y_{j}$ where $X_{j},Y_{i}\sim\Gamma(1/(n-t),b)$. Consequently, the entries of $\sum_{i\in[m]}\mathbf{S}_{i}\mathbf{G}_{i}$ are a sum of $m\cdot\lceil\frac{n}{n-t}\rceil$ matrices where each entry is either 00 or a Laplacian distributed random variable with mean 00 and scale $b$. Using Lemma 5.3, we therefore have with probability $1-\beta$.
$\frac{1}{m}\left\lVert\sum_{i\in[m]}\mathbf{S}_{i}\mathbf{G}_{i}\right\rVert_{F}^{2}$
$\in O(1)\cdot m^{2}\cdot\frac{n}{n-t}db^{2}\log^{2}\frac{m\cdot\frac{n}{n-t}}{\beta}$
$\in O(1)\cdot m^{2}\cdot db^{2}\log^{2}\frac{m}{\beta},$
where we used $\frac{n}{n-t}\leq 2$ by assumption. Since $\mathbf{S}$ is a dense Rademacher matrix, we have $m=\frac{k+\log\beta^{-1}}{\alpha_{\mathbf{S}}^{2}}$ (Achlioptas, 2003; Woodruff, 2014) and $s=m$, and $b=2\eta m^{2}d/\varepsilon$ following Corollary 4.10.
Thus we obtain, with probability $1-\beta$, a $(1+\alpha_{\mathbf{S}})$ multiplicative approximation using the same calculation as in Theorem 5.4 and a suitable rescaling, and an additive $O(\frac{k^{3}d^{3/2}}{\alpha_{\mathbf{S}}^{6}}\varepsilon^{-1}\log^{2}\frac{k}{\alpha_{\mathbf{S}}\beta})$ approximation. ∎
For ridge regression, we achieve the following bound.
Theorem 5.8.
Let $\varepsilon\geq 0$, $\delta\in(0,1)$, $t^{\prime}<n$ and $\sigma$ be chosen as described in Theorem 4.7. Let $\mathbf{S}\sim\mathcal{D}_{m,n}^{\mathtt{sketch}}$. Define $A_{\lambda}$ that performs linear regression on the sketched noisy inputs, outputting $\text{argmin}_{\mathbf{x}}\lVert\mathbf{A}\mathbf{x}-\mathbf{b}\rVert^{2}+\lambda\lVert\mathbf{x}\rVert^{2}$, where the error is measured by $\text{min}_{\mathbf{x}}\lVert\mathbf{A}\mathbf{x}-\mathbf{b}\rVert^{2}+\lambda\lVert\mathbf{x}\rVert^{2}$. An instantiation of the LTM $(R_{D},T_{\mathbf{S}},A_{\lambda})$ specified in Equations 2 and 3, using noise distribution $D\sim\mathcal{N}(0,\sigma^{2})$ according to Corollary 4.7, which tolerates $t^{\prime}$ corrupted clients, computes linear regression parameters for any $\alpha_{\mathbf{S}}>0$ with probability $1-\beta$ with multiplicative error $1+\tilde{O}\left(\alpha_{S}+p+p^{2}\right)$ and additive error $\tilde{O}\left(p+p^{2}\right)$, where $p=\frac{d^{3}\alpha_{S}^{-5}\log^{5}\frac{1}{\beta}\cdot\varepsilon^{-2}\log\frac{1}{\delta}}{\lambda}$.
Proof.
@@
Let $\varepsilon$, $\delta$, $R_{D}$ and $T_{\mathbf{S}}$ be given as in the theorem. We consider the output $A_{reg}(T_{\mathbf{S}}(R_{D}(\mathbf{a}_{1}),\dots,R_{D}(\mathbf{a}_{n})))$.
@@
If $n\leq 8m\ln(dm/\delta)+t^{\prime}$, then the $R_{D}$ outputs $0^{d}$, then the optimal $\mathbf{x}^{\prime}=0^{d}$, leading to error $\lVert\mathbf{A}\mathbf{x}^{\prime}-\mathbf{b}\rVert_{2}^{2}+\lambda\lVert\mathbf{x}^{\prime}\rVert_{2}^{2}=\lVert\mathbf{b}\rVert^{2}\leq\eta^{\prime}n\leq\eta^{\prime}m\ln(dm/\delta)+t^{\prime}$.
@@
Otherwise, $\mathcal{M}(\mathbf{A})=T_{\mathbf{S}}(R_{D}(\mathbf{a}_{1}),\dots,R_{D}(\mathbf{a}_{n}))=\frac{1}{\sqrt{s}}\sum_{i\in[s]}\mathbf{S}_{i}(\mathbf{A}+\mathbf{G}_{i})$ where $\mathbf{a}_{i}$ denotes row $i$ of $\mathbf{A}\in\mathbb{R}^{n\times d}$ and $\mathbf{G}\leftarrow\mathcal{N}(0,\sigma^{2})^{n\times d}$, is $(\varepsilon,\delta)$ differentially private.
@@
for either $\mathbf{x}=\mathbf{x}_{\text{OPT}}$ or $\mathbf{x}=\mathbf{x}^{\prime}$ and for a sufficiently large constant $\eta.$ Then we have using Lemma 5.6
$\leq(1+\alpha_{S})\|\mathbf{S}(\mathbf{A}\mathbf{x}^{\prime}-\mathbf{b})\|^{2}+\lambda\|x^{\prime}\|^{2}$
@@
$=$ $\lambda\|x^{\prime}\|^{2}+(1+\alpha_{S})\frac{1}{s}$
$\cdot\|\sum_{i\in[s]}\mathbf{S}_{i}((\mathbf{A}+\mathbf{G}_{i}-\mathbf{G}_{i})\mathbf{x}^{\prime}-\mathbf{b}+\mathbf{g}_{i}-\mathbf{g}_{i})\|^{2}$
@@
(6) $\leq$ $(1+\alpha_{S})^{2}\cdot\Bigg{(}\frac{1}{s}\|\sum_{i\in[s]}S_{i}((\mathbf{A}+\mathbf{G}_{i})\mathbf{x}^{\prime}-(\mathbf{b}+\mathbf{g}_{i})\|^{2}$
(7) $+\lambda\cdot\|\mathbf{x}^{\prime}\|^{2}\Bigg{)}+2\cdot\left(1+\frac{1}{\alpha_{S}}\right)\eta\sigma^{2}nd\log\frac{1}{\beta}(\|\mathbf{x}^{\prime}\|^{2}+1)$
@@
$\|\mathbf{x}^{\prime}\|\leq\frac{1}{\lambda}\cdot\Bigg{(}$ $\frac{1}{s}\|\sum_{i\in[s]}S_{i}((\mathbf{A}+\mathbf{G}_{i})\mathbf{x}_{\text{OPT}}-(\mathbf{b}+\mathbf{g}_{i}))\|^{2}+\lambda\cdot\|\mathbf{x}_{\text{OPT}}\|^{2}\Bigg{)}.$